Find four consecutive odd integers such that twice the sum of the first and third is 17 greater than the fourth.
Answer by ewatrrr(24785)

four consecutive odd integers: x,x+2,x+4,x+6
.......
2(2x+4) = (x+6) + 17
4x + 8 = x + 23
3x = 15
x = 5
.......
The four consecutive odd integers are: 5, 7, 9, 11
And...checking
2•14 = 11 + 17 = 28